OpenStudy (anonymous):

combining like terms the problem is 3t-18t+4=-26 @mitchal

OpenStudy (anonymous):

alright so 3t muis 18t is what?

OpenStudy (anonymous):

15t

OpenStudy (anonymous):

not quite, 3t muis 18t becomes a negative

OpenStudy (anonymous):

3 muis 18 is -15 ;)

OpenStudy (anonymous):

okay and then lol

OpenStudy (anonymous):

okay so it -15t+4=-24 and youre done

OpenStudy (anonymous):

you cant combine like terms on the +4 and -24 cause they are on other side of the equation

OpenStudy (anonymous):

okay thx boo

DivineSolar (divinesolar):

-24 is wrong :)

OpenStudy (anonymous):

-26 i meant

DivineSolar (divinesolar):

There you go. -15t + 4 = -26

DivineSolar (divinesolar):

Thats only the first step simplification.

OpenStudy (anonymous):

sorry my finger is typing to fast than my brain XD

DivineSolar (divinesolar):

Then you subtract 4 from both sides. which looks like -> -15t + 4 - 4 = -26 - 4, Which makes what shay?

OpenStudy (anonymous):

umm tht would be -22?

DivineSolar (divinesolar):

No Negative - a negative is addition.

OpenStudy (anonymous):

so 28

DivineSolar (divinesolar):

So -26 - 4 makes?

OpenStudy (anonymous):

okay you start on the -26 and you go 4 more left, where would it stop?

DivineSolar (divinesolar):

It should not be hard, it would be -15t = -??

DivineSolar (divinesolar):

-26 - 4 put that in a calculator if you cannot solve it yourself :)

OpenStudy (anonymous):

it would be 28

OpenStudy (anonymous):

like the other eqaution i solve it should be 16x=10

OpenStudy (anonymous):

no no darling it is -30

DivineSolar (divinesolar):

How do you get 28 girl thats a 2 that makes 28 not a 4. :)

DivineSolar (divinesolar):

so that gives you -15t = -30

DivineSolar (divinesolar):

Now we divide both sides by -15.

OpenStudy (anonymous):

no we cant solve any more divinesolar

DivineSolar (divinesolar):

Which looks like -> \[\frac{ -15t }{ -15} = \frac{ -30 }{ -15 }\]

OpenStudy (anonymous):

she only wants to know how to combine like terms, not "solve the equation"

DivineSolar (divinesolar):

Yeah you can pal.

DivineSolar (divinesolar):

Ah ok well if you solve it you get t = 2 there is an extra bit.

OpenStudy (anonymous):

lol okay @DivineSolar

OpenStudy (anonymous):

i mean you would be right @DivineSolar if we were to solve it but like i said, were only combining like terms

DivineSolar (divinesolar):

Yeah i know. Just realized was just combining :)!

OpenStudy (anonymous):

so the answer shay would be -15t+4=-26

OpenStudy (anonymous):

thx guys @mitchal and @DivineSolar

DivineSolar (divinesolar):

Sure.

OpenStudy (anonymous):

it okay @DivineSolar you did good though ;) I give you a best response for it anyways